OpenStudy (turingtest):

on behalf of @jrzyby1 solve the system x+4y-z=3 (1) x+5y-2z=5 (2) 3x+12y-3z=9 (3)

OpenStudy (lgbasallote):

x + 4y - z = 3 x + 5y - 2z = 5 subtract the equations 9y + z = -2 <-- (4)

OpenStudy (unklerhaukus):

times the first equation by three, then find the difference with the third equation

OpenStudy (unklerhaukus):

(1),(3) are the same

OpenStudy (unklerhaukus):

infinite solutions

OpenStudy (anonymous):

Since 1 and 3 are the same, aren't there an infinite number of solutions?

OpenStudy (lgbasallote):

hmm yep equation 1 and 3 wont intersect

OpenStudy (anonymous):

gotcha.

OpenStudy (unklerhaukus):

there is an infinite line of solutions

OpenStudy (anonymous):

I wass't sure and I tried to put the second equation into the first to see if it could be worked, hence the origin of all this. lol

OpenStudy (anonymous):

cool. thank you much

OpenStudy (turingtest):

@jrzyby1 fyi to ask a new question you will have to close your old post (which I think you left open)